OpenStudy (anonymous):

simplify square root of 3 times square root of 12 minus square root of x

OpenStudy (anonymous):

Hey

OpenStudy (anonymous):

\[\sqrt{3}(\sqrt{12}-\sqrt{x})\]

OpenStudy (anonymous):

ok so you expand the bracket like so... √(3)√(12) - √(3)√(x) Now √(a)√(b)=√(ab) so √(3)√(12)=√(3 x 12) = √(36)=6 and √(3)√(x)=√(3x) so we have 6-√(3x)

OpenStudy (anonymous):

that ok?